OpenStudy (anonymous):

raman has 2 daughters deepa and anju. present age of ramanin years is nine more than that of twice the sum of deepa and anju. five years hence the age of raman will be or more than one and half times the sumof ages of deepa and anju. determine the age of raman.

OpenStudy (anonymous):

let raman age be x and sum of age of deepa and anju be y then, x = 2(y+z) + 9....i (x+5) = 3/2(y+10) ...ii Solve for x

OpenStudy (anonymous):

hey saurav.... hav 2 Qs abt this answer...what is Z and can u please explain the 2nd equation?